Find the average rate of change of the function from to f(x)= square root of x from x1=4 to x2=121
1
Expert's answer
2020-09-14T18:45:13-0400

"f(x) = \sqrt x"

"x_1=4, x_2 = 121"

Main formula: "A(x) = \frac{\Delta y}{ \Delta x} = \frac{f(x+h) -f(x)}{h}"

Find "f(x+h)" :

"f(x+h) = \sqrt {x+h}"


Find "h":

"h = x_2 - x_1 = 121 - 4 = 117"


Find the average rate of change by using our main formula:


"A(x) = \frac{f(x+h) - f(x)}{h} = \frac{\sqrt{4+117} - \sqrt{4}}{117} = \frac{11 - 2}{117} = \frac{9}{117} = \frac{1}{3}"


The answer is: "\bold{\frac{1}{3}}"